# Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v1

@@ A.1 Computational Cost Analysis
Method NMI values at different wall-clock training times NMI @ 10 min NMI @ 20 min NMI @ 30 min NMI @ 40 min NMI @ 50 min NMI @ 60 min VAE $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ IAF-VAE $0.54$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AAE $0.61$ $0.57$ $0.57$ $0.57$ $0.57$ $0.57$ DiffVAE (T=5) warm up $0.63$ $0.63$ $0.66$ $0.66$ $0.66$ DiffVAE (T=10) warm up $\mathbf{0.64}$ $\mathbf{0.68}$ $\mathbf{0.70}$ $\mathbf{0.70}$ $\mathbf{0.70}$ DiffVAE (T=20) warm up $0.50$ $0.51$ $0.56$ $0.64$ $0.68$ DiffVAE (T=50) warm up $0.52$ $0.54$ $0.51$ $0.59$ $0.59$
@@
We conduct a computational cost analysis between the baselines and DD-VAE with various timesteps on the genotype clustering/visualization experiments. The table Table 4 below shows that DD-VAE outperforms baselines at all timestamps and continues to improve after the baselines have plateaued.
A.2 Priors
@@
Pinwheel. This distribution was used in (Johnson et al., 2016). We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in Marsland (2014). We add noise $\sigma=0.1$ to the prior. For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior, before we add noise to it, can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
@@
VAE and IAF-VAE requires that we can evaluate the prior density. To do this, for all priors, we evaluate the density by fitting a kernel density estimator with mixture of gaussian kernel with bandwidth equal to 0.005, 0.008, 0.01, 0.03, and 0.05.
A.3 Model Architecture
@@
All methods use the same architecture for encoder $q_{\bm{\phi}}(z|x)$ and decoder $p_{\bm{\theta}}(x|z)$, excluding the extra parts specific to each method which we describe below, for the same dataset. For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units. For CIFAR-10, the encoder is a 4-layer convolutional neural network with (16, 32, 64, 128) channels with a linear layer on top, and the decoder is a 4-layer tranposed convolutional neural network with (64, 32, 16, 3) channels where a linear layer is used to first turn the feature dimension from 2 to 64.
@@
IAF-VAE employs 4 IAF transformations on top of the encoder, each is implemented with a 4-layer MADE. The number of hidden units in MADE is 24. The ordering is reversed between every other IAF transformation.
@@
AAE has a discriminator, used in adversarial training, which is a multi-layer perceptron with two hidden layers, each with 1000 hidden units.
@@
DD-VAE has a diffusion model on top of the encoder. The time-conditioned reverse diffusion distribution is implemented with a 5-layer time-conditioned multi-layer perceptron, each with 128 hidden units. A time-conditioned linear layer learns an additional embedding for each timestep and adds it to the output of the linear layer.
A.4 Training Details
For training, we use Adam optimizer and latent size of 2 for all of our experiments. The training details of each algorithm are detailed below:
@@ A.4 Training Details
VAE. The batch size is set to 128. The number of epochs is 200 for unsupervised and clustering experiments and 50 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For semi-supervised MNIST experiments, the kl divergence weight is set to be 0.01, while for semi-supervised CIFAR-10 experiments, the kl divergence weight is set to be 0.01. For other experiments, the KL divergence weight is set with a schedule linear on number of epochs going from 0 to 0.01. We also have a weight of 5 multiplied to the prior density.
@@
IAF-VAE. The batch size, number of epochs, learning rate, loss, KL divergence weight, and prior density weight are the same as VAE. The context size, i.e., the size of features used to initialize the flow layers for different datat point, is 10.
@@
AAE. The batch size is set to 128. The number of epochs is 200 for all experiments. The learning rate is 0.0002. The loss is MSE for all experiments. To stabilize the training, we add noise to the input to the discriminator with sigma 0.3 at the start and lower it by 0.1 for every 50 epochs. The noise equals to 0 at epoch 150.
@@
DD-VAE. The batch size is set to 128 for most experiments, except for semi-supervied experiments where the batch size is 1024. The number of epochs is 200 for unsupervised and clustering experiments and 30 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For unsupervised MNIST and CIFAR-10 experiments, the KL divergence weight is set to 0.003. For semi-supervised MNIST experiment, we use KL divergence weight of 0.1. For semi-supervised CIFAR-10 experiment, we use KL divergence weight of 0.5. For clustering experiment, we use KL divergence weight of 0.005. The number of timesteps is 20 for unsupervised and clustering experiments and 100 for semi-supervised experiments.
A.5 Genotype Analysis Experiments Details
@@
Method Pinwheel Swiss Roll Square Acc Latent NLL Acc Latent NLL Acc Latent NLL VAE $0.16\pm 0.01$ $86.9\pm 99.03$ $0.15\pm 0.01$ $360.41\pm 185.87$ $0.15\pm 0.01$ $63.3\pm 70.64$ IAF-VAE $0.19\pm 0.01$ $3.12\pm 1.16$ $0.16\pm 0.01$ $189.39\pm 57.48$ $0.18\pm 0.01$ $0.99\pm 0.11$ AAE $0.23\pm 0.00$ $2.30\pm 0.07$ $\bm{0.23\pm 0.01}$ $\bm{2.84\pm 0.08}$ $0.20\pm 0.01$ $1.46\pm 0.49$ HVAE w/ flow $0.20\pm 0.01$ $2.06\pm 0.13$ $0.20\pm 0.01$ $106.15\pm 15.30$ $0.19\pm 0.01$ $1.05\pm 0.06$ DiffVAE $\bm{0.23\pm 0.01}$ $\bm{1.37\pm 0.02}$ $\bm{0.23\pm 0.01}$ $\bm{2.80\pm 0.11}$ $\bm{0.24\pm 0.01}$ $\bm{0.86\pm 0.05}$
Table 5: Unsupervised learning on CIFAR-10. We report accuracy using KNN (K=20) classifier and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
@@
Method Pinwheel Swiss Roll Square Latent NLL Avg Class Latent NLL Avg Class Latent NLL Avg Class Latent NLL Latent NLL Latent NLL VAE $7.67\pm 1.34$ $9.84\pm 2.87$ $49.64\pm 28.12$ $128.46\pm 47.08$ $1.51\pm 0.23$ $0.65\pm 0.37$ IAF-VAE $4.54\pm 0.59$ $4.46\pm 0.99$ $64.27\pm 73.87$ $102.16\pm 106.36$ $1.04\pm 0.07$ $0.07\pm 0.09$ AAE $2.33\pm 0.02$ $2.42\pm 0.06$ $3.76\pm 0.04$ $4.23\pm 0.13$ $1.45\pm 0.12$ $1.31\pm 0.20$ DiffVAE $\bm{1.49\pm 0.09}$ $\bm{0.35\pm 0.18}$ $\bm{3.35\pm 0.27}$ $\bm{2.28\pm 0.39}$ $\bm{0.97\pm 0.03}$ $\bm{-0.29\pm 0.04}$
@@
Table 6: Semi-supervised learning on CIFAR-10 (10,000 labels). We report average class latent negative log-likelihood and latent negative log-likelihood with pinwheel, swiss roll, and square priors.
@@
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square VAE $6.18\pm 0.93$ $82.96\pm 99.30$ $25.07\pm 62.41$ IAF-VAE $2.06\pm 0.21$ $12.13\pm 18.28$ $1.53\pm 0.82$ AAE $1.94\pm 0.09$ $3.43\pm 0.14$ $1.97\pm 0.70$ HVAE w/ flow $2.18\pm 0.13$ $4.17\pm 1.23$ $\mathbf{1.25\pm 0.08}$ DiffVAE $\mathbf{1.57\pm 0.10}$ $\mathbf{3.19\pm 0.07}$ $\mathbf{1.24\pm 0.13}$ VAE with Diffusion (standard ELBO) $6.09$ $510.53$ $265.25$
@@
Table 7: Unsupervised learning on MNIST, including the failed results of VAE with Diffusion
Model Training Objective Approximating Family Sample-based Prior Auxiliary Variable Tasks Simplified Graphical Illustration VAE ELBO Diagonal Gaussian No No Density estimation $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ IAF-VAE ELBO Normalizing flow No Yes Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{x}$ AAE Adversarial training Adversarial generator Yes No Visualization $\mathbf{x}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ HVAE w/o and w/ flow ELBO Factorial Normal / Normalizing flow No Yes Density estimation / High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ ADGM ELBO Non-Gaussian No Yes Density estimation $\mathbf{x}\rightarrow\mathbf{a}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$ LDM ELBO Diagonal Gaussian No Yes High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ LSGM ELBO+score matching Diagonal Gaussian No Yes High-quality sample generation $\mathbf{x}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{z}_{T}\rightarrow\mathbf{z}_{0}\rightarrow\mathbf{x}$ DiffVAE ELBO+sleep term Denoising diffusion Yes Yes Density estimation / Visualization $\mathbf{x}\rightarrow\mathbf{z}_{T}(\mathbf{y})\rightarrow\mathbf{z}_{0}(\mathbf{z})\rightarrow\mathbf{x}$
Table 8: Comparison of DiffVAE to other relevant methods

### v3

@@ Appendix D Computational Cost Analysis
Method NMI values at different wall-clock training times NMI @ 10 min NMI @ 20 min NMI @ 30 min NMI @ 40 min NMI @ 50 min NMI @ 60 min AEVB $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AEVB-IAF $0.54$ $0.52$ $0.52$ $0.52$ $0.52$ $0.52$ AAEB $0.61$ $0.57$ $0.57$ $0.57$ $0.57$ $0.57$ DDVI (T=5) warm up $0.63$ $0.63$ $0.66$ $0.66$ $0.66$ DDVI (T=10) warm up $\mathbf{0.64}$ $\mathbf{0.68}$ $\mathbf{0.70}$ $\mathbf{0.70}$ $\mathbf{0.70}$ DDVI (T=20) warm up $0.50$ $0.51$ $0.56$ $0.64$ $0.68$ DDVI (T=50) warm up $0.52$ $0.54$ $0.51$ $0.59$ $0.59$
@@
We conduct a computational cost analysis between the baselines and DDVI with various timesteps on the genotype clustering/visualization experiments. Table 4 shows that DDVI outperforms baselines at all timestamps and continues to improve after the baselines have plateaued.
Appendix E Connections to Diffusion Samplers
Diffusion sampling [3, 46, 41, 45, 29, 32, 1] mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) [46], denoising diffusion sampler (DDS) [41], and time-reversed diffusion sampler (DIS) [3], which have been unified by Richter et al. [29]. Some other works [45, 1] use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. Sendera et al. [32] benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps. Overall, there are indeed some strong connections between these works and ours:
  • They also focus on variational methods that directly fit a parametric family of tractable distributions (given by controlled SDEs) to the target density.
  • They cast the density estimation/sampling problem into an optimization problem over a control objective, which learns control drifts (and diffusion) parameterized by neural networks.
But we would like to clarify that there are also some clear differences between them:
  • The diffusion-structured samplers only focus on density estimation/sampling but ignore the problem of learning a generative model, which is one of the main focuses of our work. We aim to perform more accurate variational inference using an auxiliary variable model augmented by diffusion models to improve generative modeling. In our setting, $p_{\bm{\theta}}(z|x)$ is a moving target density, as we jointly learn ${\bm{\theta}}$ with ${\bm{\phi}}$, as opposed to a static target density that diffusion-structured samplers are designed to solve.
  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy [46] or off-policy [21] trajectory-based optimization. Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
  • In Equation 9, we are minimizing the forward KL divergence $\mathcal{D}_{KL}(p_{\bm{\theta}}||q_{\bm{\phi}})$, where diffusion samplers are minimizing the reverse $\mathcal{D}_{KL}(q_{\bm{\phi}}||p_{\bm{\theta}})$.
We also summarize the connections and differences in the table below.
Diffusion Samplers GFlowNet-based DDVI (ours)
[3, 46, 41, 1] [45, 29, 32, 21]
Tasks Sampling, density estimation Sampling, density estimation Learning, Sampling, Dimensionality reduction
Model Family for $p$ Any energy-based Any energy-based Latent with tractable $p(x|z),p(z)$
Model Family for $q$ Markov chain Markov chain Markov chain
Objective $D_{KL}(q||p)$ with regularizer Trajectory balance objective $D_{KL}(p||q)$ with ELBO
Algorithm Gradient descent (with reference process), importance sampling RL-motivated off-policy optimization (replay buffers, Thompson sampling, etc.) Gradient descent with wake-sleep
Compatible Models Anything energy-based Anything energy-based LDA, deep latent-variable models
Applications Sampling from physics-based models, model selection based on NLL Sampling from physics-based models, model selection based on NLL Probabilistic programming, visualization
Table 5: Comparison of Diffusion-structured Samplers, GFlowNet-based Approaches, and DDVI
Appendix F Priors
@@
Pinwheel. This distribution was used in [13]. We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in [22]. For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
@@
AEVB and AEVB-IAF requires that we can evaluate the prior density. To do this, for all priors, we evaluate the density by fitting a kernel density estimator with mixture of gaussian kernel with bandwidth equal to 0.005, 0.008, 0.01, 0.03, and 0.05.
Appendix G Model Architecture
@@
All methods use the same architecture for encoder $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ and decoder $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$, excluding the extra parts specific to each method which we describe below, for the same dataset. For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units. For CIFAR-10, the encoder is a 4-layer convolutional neural network with (16, 32, 64, 128) channels with a linear layer on top, and the decoder is a 4-layer tranposed convolutional neural network with (64, 32, 16, 3) channels where a linear layer is used to first turn the feature dimension from 2 to 64.
@@
AEVB-IAF employs 4 IAF transformations on top of the encoder, each is implemented with a 4-layer MADE. The number of hidden units in MADE is 128. The ordering is reversed between every other IAF transformation.
@@
AAEB has a discriminator, used in adversarial training, which is a multi-layer perceptron with two hidden layers, each with 1000 hidden units.
@@
DDVI has a diffusion model on top of the encoder. The time-conditioned reverse diffusion distribution is implemented with a 5-layer time-conditioned multi-layer perceptron, each with 128 hidden units. A time-conditioned linear layer learns an additional embedding for each timestep and adds it to the output of the linear layer.
Appendix H Training Details
For training, we update the parameters for each batch of inputs by alternating between the ELBO phase (optimizing ${\bm{\theta}}$ and ${\bm{\phi}}$ with respect to the ELBO, i.e., the reconstruction term and the prior matching term) and the sleep phase (optimizing ${\bm{\phi}}$ with respect to the sleep term). We use Adam optimizer and latent size of 2 for all of our experiments. Each algorithm takes roughly 2 hours on a single Nvidia GeForce RTX 3090 to complete one run of experiment. The training details of each algorithm are detailed below:
@@ Appendix H Training Details
AEVB. The batch size is set to 128. The number of epochs is 200 for unsupervised and clustering experiments and 50 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For semi-supervised MNIST experiments, the kl divergence weight is set to be 0.01, while for semi-supervised CIFAR-10 experiments, the kl divergence weight is set to be 0.01. For other experiments, the KL divergence weight is set with a schedule linear on number of epochs going from 0 to 0.01. We also have a weight of 5 multiplied to the prior density.
@@
AEVB-IAF. The batch size, number of epochs, learning rate, loss, KL divergence weight, and prior density weight are the same as VAE. The context size, i.e., the size of features used to initialize the flow layers for different datat point, is 10.
@@
AAEB. The batch size is set to 128. The number of epochs is 200 for all experiments. The learning rate is 0.0002. The loss is MSE for all experiments. To stabilize the training, we add noise to the input to the discriminator with sigma 0.3 at the start and lower it by 0.1 for every 50 epochs. The noise equals to 0 at epoch 150.
@@
DDVI. The batch size is set to 128 for most experiments, except for semi-supervied experiments where the batch size is 1024. The number of epochs is 200 for unsupervised and clustering experiments and 30 for semi-supervised experiments. The learning rate is 0.0001. The loss is BCE for MNIST and CIFAR-10 experiments and MSE for genotype analysis experiments. For unsupervised MNIST and CIFAR-10 experiments, the KL divergence weight is set to 0.003. For semi-supervised MNIST experiment, we use KL divergence weight of 0.1. For semi-supervised CIFAR-10 experiment, we use KL divergence weight of 0.5. For clustering experiment, we use KL divergence weight of 0.005. The number of timesteps is 20 for unsupervised and clustering experiments and 100 for semi-supervised experiments.
Appendix I Genotype Analysis Experiments Details
@@
Method Pinwheel Swiss Roll Square ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$ AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\bm{-2.10\pm 0.31}$ $1.71\pm 0.77$ AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\bm{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$ H-AEVB $-4.42\pm 0.46$ $\bm{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$ PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$ DDVI $\bm{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\bm{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\bm{1.58\pm 0.09}$
@@
Table 6: Unsupervised learning on CIFAR-10. We report ELBO and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
@@
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\bm{-1.60\pm 0.29}$ $\bm{0.49\pm 0.01}$ $\bm{1.09\pm 0.05}$ $\bm{-4.13\pm 1.51}$ $\bm{0.47\pm 0.09}$ $\bm{2.29\pm 0.08}$ $\bm{-1.73\pm 0.64}$ $\bm{0.49\pm 0.01}$ $\bm{1.48\pm 0.02}$
Table 7: Semi-supervised learning on CIFAR-10 (10,000 labels). We report ELBO, accuracy using KNN (K=20) classifier (Acc), and latent negative log-likelihood (Latent NLL) with pinwheel, swiss roll, and square priors.
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\bm{1.27\pm 0.21}$ $\bm{3.86\pm 1.17}$ $\bm{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
@@
Table 8: Unsupervised learning on MNIST, including the results of DDVI without the sleep term.
Appendix J Data Licenses
Yann LeCun and Corinna Cortes hold the copyright of MNIST dataset, which is a derivative work from original NIST datasets. CIFAR-10 is under MIT license. 1000 Genomes dataset is under the creative commons Attribution-NonCommercial-ShareAlike 3.0 Unported licence.